LASER SPECTROSCOPY OF THE Ã ^{2}Π − X̃ ^{2}Σ^{+} AND C̃ ^{2}Π − Ã ^{2}Π TRANSITIONS OF SrOD AND THE Ã ^{2}Π − X̃ ^{2}Σ^{+} TRANSITION OF BaOH

\underline{SHANSHAN YU}, JIN-GUO WANG, PHILLIP M. SHERIDAN, Department of Chemistry, University of Waterloo, Waterloo, Ontario, Canada N2L 3G1; MICHAEL J. DICK, Department of Physics, University of Waterloo, Waterloo, Ontario, Canada N2L 3G1; PETER F. BERNATH, Department of Chemistry, University of Waterloo, Waterloo, Ontario, Canada N2L 3G1.

Laser excitation spectroscopy of SrOD and BaOH and optical optical double resonance spectroscopy of SrOD have been carried out at high-resolution. SrOD and BaOH were produced by the reaction of metal vapor and D_{2}O/H_{2}O vapor in a Broida-type oven and in a laser ablation/molecular beam spectrometer, respectively. For SrOD, the Ã ^{2}Π − X̃ ^{2}Σ^{+} and C̃ ^{2}Π−Ã ^{2}Π transitions were observed and rotationally analyzed for the first time. Combined with the previous microwave data from the literature, the present data were fitted using the usual ^{2}Π and ^{2}Σ Hamiltonians, and spectroscopic constants were obtained for the X̃ ^{2}Σ^{+} (000), Ã ^{2}Π (000) and C̃ ^{2}Π (000) states of SrOD. For BaOH, the Ã ^{2}Π − X̃ ^{2}Σ^{+} transition was observed and rotationally assigned. Unfortunately, one spin component of this transition could not be calibrated since our Ti:Sapphire laser had very low output power.